OpenStudy (anonymous):

Rewrite the expression ln(x-3) + ln (x+3) - ½ ln(x²+9) as a single logarithm

OpenStudy (anonymous):

\[\ln(x-3) + \ln (x+3) - ½ \ln(x²+9)=\ln{(x-3)(x+3) \over \sqrt{x^2+9}}=\ln{x^2-9 \over \sqrt{x^2+9}}\]

OpenStudy (anonymous):

Thank you ! :)